OpenStudy (anonymous):

Suppose that you invest $100 today in a risk-free investment and let the 5 percent annual interest rate compound. Rounded to full dollars, what will be the value of your investment 6 years from now?

OpenStudy (anonymous):

The formula for compound interest is given as: C.I. = Value of Investment * ((1+ (I.R./100))^n) C.I. is Compound Interest I.R. is the Interest Rate n is the number of years Value of Investment= $100 (1+ (I.R./100)) = 1+ 0.05 =1.05 n=6 Put that into the formula and its done!

OpenStudy (anonymous):